OpenStudy (anonymous):

Rex is 30 yrs.younger than his father.his mother is three times as old as her son.what is Rex's age if his father is two years older than his mother?

OpenStudy (anonymous):

let Rex age be x his fathers age is (x+30) and his mothers age is (x+30)-2=x+28 but (x+28)=3x or 2x=28 or x=14 hence Rex is 14 yrs old

OpenStudy (anonymous):

thanks.